OpenStudy (anonymous):

ln(x^2-3x-1)-ln(x+4)=0 solve for x

myininaya (myininaya):

ln[(x^2-3x-1)/(x+3)]=0 (x^2-3x-1)/(x+3)=1

myininaya (myininaya):

that 3 should be a four lol down below

OpenStudy (anonymous):

x^2-3x-1=x+4 x^2-4x-5=0 (x-5)(x+1)=0 x=5,-1

OpenStudy (anonymous):

lol k

OpenStudy (anonymous):

how did you get one on the other side of the equal sign?

OpenStudy (anonymous):

ln(x^2-3x-1)=ln(x+4) hence i got...

myininaya (myininaya):

e^blah=e^blah but if the blah is 0 then e^0=1

myininaya (myininaya):

is that too blah for you?

OpenStudy (anonymous):

ohhh ok

OpenStudy (anonymous):

I got it. :) thanks

myininaya (myininaya):

cool!